### v3

# Robust Counterfactual Explanations under Model Multiplicity Using Multi-Objective Optimization

Abstract
In recent years, explainability in machine learning has gained importance. In this context, counterfactual explanation (CE), which is an explanation method that uses examples, has attracted attention. However, it has been pointed out that CE is not robust when there are multiple machine-learning models with similar accuracy. These problems are important when using machine learning to make safe decisions. In this paper, we propose robust CEs that introduce a new viewpoint—Pareto improvement—and a method that uses multi-objective optimization to generate it. To evaluate the proposed method, we conducted experiments using both simulated and real data.The results demonstrate that the proposed method is both robust and practical. This study highlights the potential of ensuring robustness in decision-making by applying the concept of social welfare. We believe that this research can serve as a valuable foundation for various fields, including explainability in machine learning, decision-making, and action planning based on machine learning.
Keywords: counterfactual explanation, robustness, model multiplicity, pareto improvement, multi-objective optimization
1 Introduction
Artificial intelligence (AI), including machine learning, is used in many domains. However, although many machine-learning methods have high prediction accuracy, they are often considered ’black boxes’ because the processes involved are unclear owing to their complex combination of nonlinearities and interactions. Explainable AI or interpretable machine learning has become an important issue in addressing these problems [18, 1, 7]. Several such methods are available. One such method is white-box machine learning. There are also methods for ensuring the interpretability of black-box machine learning. They examine which variables are important in the overall data and which variables are important in individual data. Among these methods, one is called the counterfactual explanation (CE) [14, 10, 27].
CEs are outputs that indicate that, for a trained supervised machine-learning model, the minimum changes to the original data (explanatory variables) are needed to achieve a particular desired predictive outcome. This clarifies the factors influencing the forecast and improves the explainability of the model. For example, if a person is denied a loan by machine learning owing to a feature, the CE will suggest which features should be changed (e.g., annual income) to be approved for a loan. This method is important because it can provide suggestions for machine learning users regarding the actions they should take. CEs are also called algorithmic recourse [14]. It has also been noted that CEs are related to adversarial examples [9, 21].
When extracting CEs, the basic condition is to make the original data and generated CEs as close as possible. In addition, various other conditions have been proposed for CEs, such as closeness to the training data/plausibility, actionability (feasibility), sparsity, diversity, and so on [10]. In addition, several extraction methods have been developed depending on the availability of access to the model and the assumptions of the model’s functions (linearity, differentiability, etc.) [27].
However, the robustness of CEs has long been problematic [27]. Robustness can be considered in various ways. Jiang et al. classified the robustness of CEs into four categories [12].
(i) Robustness against model changes [26, 21, 16]: Robust (unchanged) CEs are extracted when the model changes owing to changes in data.
(ii) Robustness against model multiplicity [22]: Robust CEs are extracted when there are multiple models with the same accuracy for the same dataset of classification tasks. Specifically, the formulation uses an aggregated model to make predictions and then a subset of models to generate CEs that differ from those predictions.
(iii) Robustness against noisy executions [23]: Robust CEs are extracted such that, even if their attributes change slightly, the predictions do not change significantly.
(iv) Robustness against input changes [25]: Robust CEs are extracted under the condition that if the predictions of two similar data points are identical, then the CEs of the data are also similar.
Regarding (ii), the existence of multiple models with similar accuracy, which is the premise of the problem, is an important issue because it is sometimes observed when comparing a large number of learners using automated machine-learning tools such as Google Cloud AutoML or PyCaret. Furthermore, in such situations, CE is an important task for social applications because it can be used to identify important variables without selecting a model and to ensure safety when making further decisions. For example, it is important to select the most effective CEs when making decisions based on CEs for risky issues such as medical care or issues involving huge costs such as marketing. However, the number of relevant studies is limited [22, 17, 12].
Pawelczyk et al. [22] conducted a theoretical study on the relationship between several features of CE under model multiplicity. For example, it discusses the cost of CE (the minimum amount of change required to alter the predicted result from the original input data) under model multiplicity and shows that, while the sparse method (which minimizes the L1 or L2 norm of changes to the data) has a lower cost, the data support method (which generates explanations closer to the data distribution) is more robust in handling model multiplicity. Additionally, in the work of Leofante et al. [17], it is defined such that the predicted categories were the same in all models. Specifically, model multiplicity in class classification using a homogeneous feedforward neural network (FFNN) was analyzed using mixed-integer optimization and other methods. Jiang et al. [13] proposed an algorithm for extracting CEs that can be used universally. The CEs in this study correspond to the prediction discrepancies in model multiplicity. Specifically, after generating CEs for each model, the algorithm uses the computational argumentation method to extract CEs that satisfy several requirements such as nonemptiness and majority voting. Although there have been some studies, they have been limited to the analysis of white-box models or classifications [17] or have not proposed general-purpose solutions that could be applied to cases where the target variable is continuous and constraints can be included. In addition, there is little theoretical foundation for a consistent CE in multiple models.
In game theory, welfare economics and multi-objective optimization, the concept of Pareto efficiency describes a state in which any attempt to improve at least one objective function inevitably worsens at least one other function[20]. Once a state is Pareto efficient, no further Pareto improvements—changes that benefit one or more functions without harming others—are possible. In other words, a Pareto improvement moves the previous state closer to a Pareto-efficient outcome. However, once this efficiency is reached, no further improvements can be made. As will be discussed in more detail below, this idea can be used to define a CE that improves at least one in all models; that is, it improves consistently in all models. This idea is important when choosing safe and uncontroversial solutions based on the CE, which is a costly and risky problem in society.
In this paper, we propose robust CEs for model multiplicity by introducing a new viewpoint, Pareto improvement, and multi-objective optimization. In addition, we devised a validation index and verified its robustness through the validation of simulated data and its practicality through the validation of real data. The proposed method has four key features. The first is the use of the concept of Pareto improvement to address the problem of model multiplicity; the second is that it can be used in practice to select safe solutions based on CE for risky problems; the third is that it can be applied not only to class classification but also to continuous target variables, such as in regression, and is sufficiently versatile to allow the free inclusion of constraint conditions; and fourth, by extracting a variety of CEs, it is possible to select a solution according to the user’s preference.
Leofante et al. [17] also defined a CE that was consistent among all models, similar to that in the present study. However, their analysis was limited to white-box models. Another difference is that a consistent CE is defined as a Pareto solution. Jiang et al. [13] extracted and refined CEs from their model and then refined them [13]. However, the method used in this study extracted multiple consistent CEs directly from multiple models. These differences include the flexibility to handle regression, incorporate constraints on CEs, and the possibility of extracting a variety of CEs.
Similar studies have used MOO for CEs. They primarily focus on detecting CEs to satisfy multiple conditions (e.g., Validity, Proximity, Sparsity) in the generation of CEs in one model [3]. However, this study focuses on the model multiplicity problem and uses MOO to ensure the robustness of CEs. Therefore, the problem settings, research positioning, and procedures of the proposed methods differ.
By introducing the concepts of Pareto improvement and multi-objective optimization, this research contributes to a wide range of areas, including the study of explainability in machine learning, decision-making, and action planning based on machine learning.
Section 2 presents the proposed method. Section 3 presents the validation using simulated and real data, and Section 4 provides a discussion.
2 Method
In the following sections, Section 2.1 describes the problem setup, Section 2.2 explains the multi-objective optimization as a prerequisite for the proposed method, Section 2.3 describes the proposed method, and Section 2.4 explains the evaluation method.
2.1 Problem Setting
We set up our problem as follows. We have data $\mathcal{D}$, consisting of $n$ pairs of $y_{i}$ (scalars) and $X_{i}$ ($r$-dimensional vectors), where $i$ is an index per sample. Let $y_{i}\in\mathcal{Y}\subseteq\mathbb{R}$ and $X_{i}\in\mathcal{X}\subseteq\mathbb{R}^{r}$. $\mathcal{Y}$ and $\mathcal{X}$ are feature spaces.
$\mathcal{D}=\left\{(y_{i},X_{i})\right\}_{i=1}^{n}$
There are $m$ machine-learning models $f_{j=1,\dots,m}:\mathcal{X}\to\mathcal{Y}$ estimated from $\mathcal{D}$, above a certain accuracy. Based on the above, the objective of this study is to obtain the solution $X_{cf}^{*}$ to the following problem:
$X_{cf}^{*}=\operatorname*{argmin}_{X_{c}\in\mathcal{X}}\left(\text{loss}(y_{t},f_{1}(X_{cf})),\text{loss}(y_{t},f_{2}(X_{cf})),\dots,\text{loss}(y_{t},f_{m}(X_{cf}))\right)$
subject to
$d(X_{b},X_{cf})\leq C,$
$g_{j}(X_{cf})\geq 0,\quad j=1,\dots,J,$
$h_{k}(X_{cf})=0,\quad k=1,\dots,K.$
where $y_{t}\in\mathcal{Y}$ is a target value, $X_{b}\in\mathcal{X}$ is a base data to be explained, and $X_{cf}$ is a candidate for counterfactual explanations. loss is the loss function between $y_{t}$ and $f_{j}(X_{cf})$. Specifically, when $y$ is a continuous variable, the squared or absolute error is used. When $y$ is a categorical variable, the cross-entropy error is used. The $m$ values of the losses are vectors. $d$ is a distance function that uses the Euclidean distance, squared Euclidean distance, etc. $C$ is the upper bound of the distance function. The $J$ types of functions $g_{j}:\mathcal{X}\to\mathbb{R}$ and $K$ types of functions $h_{k}:\mathcal{X}\to\mathbb{R}$ are the functions of constraints. In summary, this study seeks to find $X_{cf}$ that is less than or equal to $C$ with respect to the base $X_{b}$, satisfies certain constraints, and is better than the other solutions in the loss function for all models. Such a solution with repeated Pareto improvements is called the Pareto solution, the details of which are described in Section 2.2.
2.2 Multi-Objective Optimization
In this section, we provide a basic explanation of multi-objective optimization and Pareto solutions [6]. We introduce $L$ ($l=1,\dots,L$) objective functions $F_{l}:E\to\mathbb{R}$ corresponding to $r$-dimensional variables $\theta\in E$. Let $E$ be a domain and $E\subseteq\mathbb{R}^{r}$. Let $F:E\to\mathbb{R}^{L}$ be the function that summarizes them. We solve the following optimization problem:
$\min_{\theta\in C}F(\theta),\quad\text{where}\quad F(\theta)=\left(F_{1}(\theta),F_{2}(\theta),\dots,F_{L}(\theta)\right).$
Constraints can also be included. The Pareto solution, which is the solution to this problem, refers to a $\theta^{*}$ such that there exists no $\theta$ for which $F(\theta)\leq F(\theta^{*})$ and $F(\theta)\neq F(\theta^{*})$. This is also called a non-dominated or efficient solution. The condition $F(\theta)\leq F(\theta^{*})$ means that $F_{l}(\theta)\leq F_{l}(\theta^{*})$ for all $l$. The Pareto solution set $\Theta^{*}\ni\theta^{*}$ is also called the Pareto front in the objective function space.
There are many methods for computing Pareto solutions in multi-objective optimization. For example, there are methods that use a weighted sum of multiple objective functions and the $\epsilon$-constraint method, which optimizes a specific objective function and defines all other functions as constraints. In addition to those that seek a specific optimal solution, there are also those that seek a set of Pareto solutions, as described above [4].
Recently, several methods have been proposed for computing Pareto solution sets, including evolutionary computation and descent methods. A well-known method that uses evolutionary computation is the fast elitist non-dominated sorting genetic algorithm (NSGA-II) [8, 5]. This method identifies a set of solutions by iteratively performing ranking, selection, crossover, and mutation operations on a population of candidate solutions. During this process, the crowding distance metric is utilized to ensure a diverse spread of solutions across the Pareto front.
Because evolutionary computation is an approximate solution method (heuristic), there is no guarantee that a Pareto solution set can be obtained. Therefore, the output is often a non-dominated, non-inferior solution set among the solution sets obtained in the solution search process. However, methods using evolutionary computation have the advantage of being able to extract a wide variety of solutions, are robust to nonlinear and nonconvex functions, and eliminate the need to differentiate the objective function. However, this approach is computationally expensive. Other approaches exist that use the descent method [19]. Although the descent method converges quickly and is computationally inexpensive, it has problems such as falling into local solutions, requiring differentiability of the objective function, setting weights among the objective functions, and difficulty in obtaining diverse solutions.
2.3 Proposed Method
We describe the procedures proposed in this paper based on this setup.
1. Split $\mathcal{D}$ into training data $\mathcal{D}_{\text{train}}\subset\mathcal{D}$ and test data $\mathcal{D}_{\text{test}}\subset\mathcal{D}$ ($\mathcal{D}_{\text{train}}\cap\mathcal{D}_{\text{test}}=\emptyset$).
2. Set up $M$ models in advance. Estimate each model $f_{j}$ based on $\mathcal{D}_{\text{train}}$ and calculate the accuracy of prediction using $\mathcal{D}_{\text{test}}$.
3. Select $m$ models based on their accuracy.
4. Derive $S$ solutions $X_{cf,s}^{*}$ ($s=1,\dots,S$) for $m$ models by multi-objective optimization.
In Process 2, $M$ various models were prepared in advance. The MSE was used for accuracy because continuous variables were used for $y$ in this study. In Process 3, sorting was performed based on accuracy, and the top $m$ models were used. Another possible method is to set a threshold value and select models that exceed it. This selection may be a response to the increase in dissimilarity caused by the introduction of a Pareto solution. For Processes 1–3, it is possible to select the top $m$ models from a large number of models using automated machine learning. In Process 4, evolutionary computation is used because the machine-learning function is nonlinear, model-independent, and derives a wide variety of solutions. After Process 4, it is possible to select a solution among the $S$ solutions according to the user’s preference or to select the solution closest to all solutions (medoid, close to centroid, etc.) to select a safer solution.
2.4 Evaluation Method
In this study, we developed evaluation indices. When there were $S$ counterfactual explanations (CEs) for a base dataset, we used the average value described below to evaluate the method. When evaluating the method as a whole, we compared the mean values of the CEs for multiple base datasets.
Validity(Val): This is an evaluation of the closeness of the prediction by CE to the desired value [27]. Specifically, the value of the loss function was used. The smaller this value, the better the CE. However, since $y_{t}=\infty$ is set in this study, alternatively, a larger predicted value $y_{t}=\infty$ is a good CE.
$\textit{Val}_{j}=\frac{\sum_{s=1}^{S}|y_{t}-f_{j}(X_{cf,s}^{*})|}{S}$
Dissimilarity(Dissim): This indicator is the opposite of proximity. The smaller this indicator, the better the CE because it corresponds to the cost.
$\textit{Dissim}=\frac{\sum_{s=1}^{S}(X_{cf,s}^{*}-X_{b})^{2}}{S}$
Plausibility(Plaus): This indicator is also called ”closeness to the training data.” The index is based on [10] and others. The smaller this indicator is, the more feasible it is, which makes it a good CE. It is used in constraints and as an evaluation indicator [27].
$\textit{Plaus}=\frac{\sum_{s=1}^{S}\min_{X_{i}\in X}(X_{cf,s}^{*}-X_{i})^{2}}{S}$
True Improvement Ratio (TIR): We propose the following method as an indicator to check robustness: When the CE is intended to increase its predicted value, we compare the value of multiple CEs input to the true function $y_{s}$ with the predicted value of the multiple CEs input to the estimated function $f(X_{cf,s}^{*})$, and compute the ratio of CEs that are improving. If this value is high, then the CE is robust. However, this can only be used when the true function is known, such as in simulation data.
$\textit{TIR}=\frac{\sum_{s=1}^{S}1(f(X_{cf,s}^{*})-y_{s}>0)}{S}$
3 Experiment
In Experiment 1, the robustness of the CE extracted by the proposed method was verified using simulated data; in Experiment 2, it was applied to real data, and its practicality was discussed.
3.1 Experiment 1: Simulation Data
Here, we apply this method and other methods to the simulation data for which the true function is known and compare them. A comparison was performed on data with complex nonlinear functions.
The four models used are:
Model 1: Linear regression
Model 2: Random Forest regression with 100 trees
Model 3: LightGBM regression with 100 boosting rounds [15]
Model 4: Multilayer perceptron regression (MLP) with one layer of 100 units and ReLU activation
which are commonly used in machine learning. The following three methods of extracting CEs are compared:
  • Method 1: Multiple CEs are generated by changing the initial value of each model. Specifically, the following optimization problem is solved using nonlinear optimization (COBYLA) [24] to generate CEs:
    $X_{cf}^{**}=\operatorname*{argmin}_{X_{cf}\in X}\left(\text{loss}(y_{t},f_{1}(X_{cf}))+\lambda d(X_{b},X_{cf})\right),$
    $\text{subject to}\quad d(X_{b},X_{cf})\leq C,\quad g_{j}(X_{cf})\geq 0,\quad j=1,\dots,J,$
    $h_{k}(X_{cf})=0,\quad k=1,\dots,K.$
    where $\lambda$ is the degree of importance of $d(X_{b},X_{cf})$.
  • Method 2: Multiple CEs are generated by changing the initial values based on a stacking model (multiple regression) using the predictions of the above models. The generation method is the same as described above.This method can also be interpreted as a solution approach for multi-objective optimization using weighted linear summation of objective functions, provided that all coefficients are positive.
  • Method 3: Multiple CEs are generated using the proposed multi-objective optimization-based method. Specifically, we compare cases in which the number of models is set to 2, 3, and 4, in descending order of accuracy. For the multi-objective optimization algorithm, we adopt NSGA-II [2], which is a type of evolutionary computation described in Section 2.3. NSGA-II is suitable for this study because it can be applied to complex functions and enables diverse solutions to be obtained by utilizing the crowding distance.
The model and the results used in the simulations are described below. The model incorporates interactions and nonlinear functions as follows:
(1) Model with interactions and nonlinearity
$y_{i}=2x_{i,1}-3x_{i,2}+0.5x_{i,3}+1.5x_{i,1}x_{i,2}-2x_{i,3}x_{i,4}+\sin(x_{i,4})x_{i,5}+\begin{cases}5&\text{if}\ x_{i,1}>0\\ -5&\text{if}\ x_{i,1}\leq 0\end{cases}+\epsilon_{i}$
(2) Model with interactions and nonlinearity
$y_{i}=\sin(\pi x_{i,1}x_{i,2})+\sin(\pi x_{i,3}x_{i,4})+x_{i,5}^{2}-0.5x_{i,1}x_{i,3}^{2}+0.7x_{i,2}x_{i,4}x_{i,5}+\epsilon_{i}$
where $x_{i,1},x_{i,2},x_{i,3},x_{i,4},x_{i,5}$ are uniform random numbers in the range $[-10,10]$, and $\epsilon_{i}$ is a standard normal random variable with a mean of 0 and a variance of 1. The index $i=1,\dots,1000$ represents a sample. The mean and standard deviation for (1) are 1.733 and 84.931, respectively, whereas those for (2) are 36.262 and 188.296, respectively.
First, the accuracies of the models were compared (Table 1, Table 2).The dataset was randomly split 20 times with a training size of 0.7 and a test size of 0.3. The mean MSEs obtained from these splits are compared below.
Table 1: Accuracy of each model (case 1)
Model MSE
Model1 (Linear Regression) $170.802\pm 22.320$
Model2 (Random Forest) $182.365\pm 19.243$
Model3 (LightGBM) $188.093\pm 24.742$
Model4 (MLP) $176.841\pm 20.610$
Stacking Model $200.146\pm 28.423$
Table 2: Accuracy of each model (case 2)
Model MSE
Model1 (Linear Regression) $6741.662\pm 418.253$
Model2 (Random Forest) $2233.613\pm 287.310$
Model3 (LightGBM) $573.673\pm 81.060$
Model4 (MLP) $2082.787\pm 203.769$
Stacking Model $604.770\pm 97.909$
Next, we evaluated these methods (Table 3 ,Table 4). Specifically, we set $y_{t}=\infty$. Twenty CEs were generated for each randomly selected base. Using these multiple CEs, the evaluation metrics described in Section 2.2 were calculated. Finally, the mean values of 50 cases were compared.
Table 3: Evaluation metrics for CEs (case 1)
Table 4: Evaluation metrics for CEs (case 2)
The parameters were set as $C=3$ and $\lambda=2$. For the calculation of val, since $y_{t}=\infty$, this study simply used the predicted values of $y$, where higher values indicated better evaluations. Moreover, as val and dissim or plaus tended to increase together, the ratios of their average values, ave val/ave dissim and ave val/ave plaus, were calculated to compare their balance. For Method 3, the average of multiple ave val values was used as ave val. The top three lowest values for dissim and plaus, and the top three highest values for FIR, ave val/ave dissim and ave val/ave plaus, are highlighted in bold.
In Case 1 (Table 3), val was higher for Method 2. Although Method 3 showed lower predictions for scenarios with two or more models, improvements were evident in all cases. However, it should be noted that there was a tradeoff between dissim. For dissim, Method 3 consistently showed smaller values. Similarly, plaus was also smaller for Method 3. On the other hand, FIR was higher for Method 3 (except in the 2-model version), indicating greater robustness, which aligned with the objective of this study. It is worth noting that, despite smaller dissim and plaus values, FIR could sometimes be low. Furthermore, while ave val/ave dissim and ave val/ave plaus were not the highest, they remained relatively high even under conditions where robustness was maintained.
In Case 2 (Table 4), val was higher using Method 2. Although Method 3 showed lower predictions for scenarios with two or more models, improvements were evident in all cases. Additionally, dissim was consistently smaller for Method 3, as was plaus. On the other hand, FIR was consistently higher for Method 3. However, it should be noted that FIR could sometimes be low despite the smaller dissim and plaus values. Furthermore, ave val/ave dissim and ave val/ave plaus remained relatively high, even under conditions where robustness was maintained.
Additionally, we aimed to infer the impact of $C$ and $\lambda$.The value of $C$ increased, the range of possible CEs expanded, leading to an increase in FIR across all methods. As the value of $\lambda$ affected the balance between val and dissim, where a decrease in one typically led to an increase in the other. Regarding the ave val/ave dissim , we cannot assert a single, definitive trend because the functions may increase or decrease simultaneously, or there may be a trade-off relationship between them.
In summary, even in Case 1 and Case 2, compared to Method 1 and Method 2, Method 3 (using 2-model or 3-model)—where models were selected in order of accuracy—maintained higher ave val/ave dissim and ave val/ave plaus ratios, achieved consistently higher FIR, and thus could be confirmed as robust.
3.2 Experiment 2: Real Data
The proposed method is applied to real-world data, where the true model is unknown. This allowed us to investigate the impacts of specific variables, evaluate their practicality, and examine their potential applications in practical scenarios.
Following is an overview of the data. The purpose of this survey was to investigate the factors influencing academic achievement among Japanese high school students. The survey was conducted in February 2022. The survey was conducted on 500 subjects (males and females aged 15 to 18) throughout Japan. The survey method was an Internet survey, and the sample was collected so that the sex and age ratios matched those of the national census. The specific survey item was academic achievement (deviation value); in Japanese educational assessments, the deviation value is commonly used to standardize scores, sex, age, and intervention (19 types) (See Appendix), using academic achievement as the target variable ($\mathcal{Y}$) and other items as features (sex, age, $T_{1}$, …, $T_{19}$).
Below are the descriptive statistics of the data (Table 5).
Table 5: Descriptive statistics
The four models listed in Section 3.1 are used. Based on this setup, we describe the results of our analysis. First, the accuracy of the model is verified (Table 6). Based on these results, we used Model 1, Model 2, and Model 4.
Table 6: Accuracy of each model
Model MSE
Model1 (Linear Regression) $170.802\pm 22.320$
Model2 (Random Forest) $182.365\pm 19.243$
Model3 (LightGBM) $188.093\pm 24.742$
Model4 (MLP) $176.841\pm 20.610$
Stacking Model $200.146\pm 28.423$
[…]
Table 7: Evaluation index of CE
For val, Method 1 showed high values, while Method 3 showed an improvement, although some values were small. For dissim, Method 3 consistently had small values. Similarly, for plaus, Method 3 also had low values. Although ave val/ave dissim and ave val/ave plaus were not the highest, they were still high, even when the model’s robustness under multiple conditions was considered.
First, the average value of each CE across 50 baseline data points was calculated. This indicated an overall trend.
Figure 1: Average of averages of CEs, including base data
(vertical axis: average; error bars: standard deviation)
Figure 2: Average of averages of CEs, excluding base data
(vertical axis: average; error bars: standard deviation)
The first was the average of the sum of the base data and improvements (Figure 1). This indicated that important variables such as $T_{1}$, $T_{3}$, $T_{8}$, $T_{11}$, $T_{13}$, $T_{14}$, $T_{17}$, and $T_{18}$ were high. The following variables are important for improving academic performance: Knowledge1 and Knowledge3, record of study time, group work, presentations, individual educational support, introduction to role models, and supplementary classes outside of school.
The above values included the base data, but we only looked at the average of the improved values (Figure 2). Thus, $T_{8}$, $T_{14}$, $T_{17}$, and $T_{18}$ must be improved. In other words, recording study time, individual educational support, introductions to seniors and other role models, and supplementary classes outside of school are important.
Next, we consider one sample as an example and review the Pareto front for that case, as well as the actual CEs obtained and their averages.
Figure 3: Pareto front(red: base data; blue: CEs)
In Figure 3, objective1, objective2, and objective3 are the prediction results of model1, model2, and model4, respectively. It was confirmed that each value improved and varied.
For this one case, 20 CEs were generated. We checked the central CE (medoid, closest_to_centroid) and average CEs (Table 8).
Table 8: Medoid_CE and closest_to_centroid_CE
For this case, it is important to improve $T_{5}$, $T_{14}$, $T_{17}$, and $T_{18}$. In other words, Knowledge5, Individual educational support, Introduction to seniors and other role models, and supplementary classes outside of school are important.
Figure 4: Average of CEs
(vertical axis: average; error bars: standard deviation)
Therefore, it is important to improve $T_{2}$, $T_{4}$, $T_{8}$, $T_{12}$, $T_{13}$, $T_{14}$, $T_{18}$, and so on. In other words, Knowledge2, Knowledge4, Recording study time, Group leadership experience, Opinion presentation and presentation, Individual educational support, and Supplementary classes outside school are important.
Thus, it is possible to propose multiple improvement options for individual base data and identify variables whose effects are more robust based on their average values.
4 Discussion
In this study, we investigated robust CEs for machine learning, particularly for the problem of model multiplicity, which has become an issue in recent years. Specifically, we introduced the concept of Pareto improvement for robust CEs against model multiplicity and proposed the extraction of robust CEs using MOO. In addition, we propose a robustness index.
The experiments were conducted using simulated and real data. In the experiment with simulated data, all of the val was improved, although some of them were low, and it was clear that dissim was low, plaus was low, and FIR was high. The experiments with real data also showed val was improved, although the val were low, and both dissim and plaus were low.
In an applied case study, we detected the variables that were important as a whole, checked the Pareto front for one case study, and confirmed that all three models improved the results. The CE selection method is also addressed.
The novelty and distinctiveness of this study lie in the following points: (1) It employs the concept of Pareto improvement to tackle issues related to model multiplicity and robustness. (2) Unlike existing CE studies that deal with model multiplicity, this approach allows the target variable to be quantitative and can flexibly incorporate constraints. (3) CE can be applied when selecting safe and effective actions for expensive and risky practical problems. (4) Extracting various CEs enables the selection of solutions that align with user preferences.
This study highlights the potential to ensure robust decision-making by applying Pareto improvement—a concept from welfare economics—together with multi-objective optimization. We believe this research can serve as a valuable foundation for various fields, including explainability in machine learning, decision-making, and action planning based on machine learning.
Future work will include determining which MOO is more optimal [11] and setting the optimization hyperparameters. In addition, different results may be obtained for method1 and method2 if the optimization method is changed. In addition, MOO has the problem of increasing computation time as the number of attributes increases and when comparing different MOO algorithms, including the descent method. Finally, the effects of the obtained CEs were tested by conducting actual intervention experiments.
Another potential avenue for future research is to incorporate causality to enhance the model’s robustness [10]. Causality pertains to the invariant structure of the underlying data, which can ultimately contribute to improved robustness. We would also like to examine the applicability of concepts such as utilitarianism and Rawlsianism to the problems addressed in this study.
Acknowledgement
This study was supported by JSPS KAKENHI Grant-in-Aid for Scientific Research (C) JP20K02004.
Appendix: Types of Educational Interventions Experienced by Each Individual
T1 Knowledge1: On average, wages (monthly and lifetime earnings) are higher for college and graduate school graduates than for high school graduates.
T2 Knowledge2: Approximate differences in average wages between college graduates and high school graduates (e.g., lifetime wages of approximately 60 million yen (men) and 70 million yen (women) are higher for college graduates (Youthful Labor Statistics 2020)).
T3 Knowledge3: On average, wages (monthly and lifetime earnings) are higher for those with science backgrounds than for those with art backgrounds.
T4 Knowledge4: Approximate difference in average wages between those with science and art backgrounds (e.g., approximately 400,000 yen (men) and 600,000 yen (women) per year, higher for those with science backgrounds (JHPS, 2010)).
T5 Knowledge5: Employment opportunities and advancement rates at universities, technical schools, and other postsecondary institutions.
T6 Considering the purpose of studying.
T7 Making and executing a study schedule.
T8 Recording study time.
T9 Recording study content.
T10 Sharing study content with friends.
T11 Group study and group work.
T12 Group leadership experience.
T13 Opinion presentation and presentation.
T14 Individual educational support.
T15 Career education.
T16 Workplace tours, internships, and other experiences.
T17 Introduction to seniors and other role models.
T18 Supplementary classes outside of school.
T19 Use of online educational materials.